OpenStudy (anonymous):

Reduction of order DE: x^2y''-7xy'+16y=0 y1=x^4 I set y2 equal to v(x)x^4, plugged in and got it to just v' and v'' terms, substituted w and i have wx^6+w'x^5=0 I'm guessing this is a really easy DE to solve because its 1st order, but i know its not linear because its equal to 0.

OpenStudy (anonymous):

Use the substitution \(x=e^t\).

OpenStudy (anonymous):

This way you'll get a second order DE with constant coefficients.

OpenStudy (jamesj):

Does the question ask you to reduce order? It's a slightly painful way to solve this problem.

OpenStudy (jamesj):

Which is to say, if you start with your original equation, \[ x^2y''-7xy'+16y=0 \] If you use a trial solution/Ansatz of \( y = x^r \), you'll find the solution.

OpenStudy (anonymous):

No it didn't say linear thats just the one I first wanted to use when i saw it. I know i don't want second order, its a reduction of order problem. I'm guessing i need to get them into dw and dx, is w' the same as dw/dx? Or can I decide that's what it represents?

OpenStudy (anonymous):

@James: It worked with me. I've got \(y''(t)-8y'(t)+16y(t)=0\), where u is a function of t=ln(x). Am I right here?

OpenStudy (anonymous):

y*

OpenStudy (anonymous):

y2 is supposed to be 1

OpenStudy (anonymous):

oops looking at the answer to the next problem, no idea what y2 is

OpenStudy (anonymous):

If this is right, then the auxiliary equation is \(m^2-8m+16=0 \implies (m-4)^2=0 \implies m=4\). So we have a repeated root at m=4. So \(y(t)=c_1e^{4t}+c_2xe^{4t}\). Npw, plug t=ln x.

OpenStudy (anonymous):

Then \(y(x)=c_1e^{4 \ln{x}}+c_2\ln(x)e^{4\ln{x}}=c_1x^4+c_2\ln(x)x^4.\)

OpenStudy (anonymous):

So yeah \(x^4\) is your first solution, and \(\ln(x)x^4\) is your second solution.

OpenStudy (anonymous):

I made a typo on above, it should be \(y(t)=c_1e^{4t}+c_2te^{4t}.\) The final answer is right.

OpenStudy (anonymous):

typo above*

OpenStudy (anonymous):

thats the best answer i've heard so far...not the way we were taught to do it but i ended up with -x^4ln(x) by seperating the variables once i got it into W's. if someone could tell me if the way i did it is ok i'd really appreciate it. (dw/dx)*x^6 = -w*x^5 dw/w = -dx/x ln(w) = -ln(x) or ln(1/x) w = 1/x v = ln(x) realized while typing it i changed the log to 1/x and kept the negative. thanks guys.

OpenStudy (anonymous):

I think what you did is correct.